OpenStudy (anonymous):

Marianne, who runs an ice cream parlor, can prepare one chocolate sundae every two minutes. She earns $1.20 for every sundae she makes. Which function relates her earnings to the number of minutes she works?

OpenStudy (ankit042):

1 sundae in 2 mins earning on 1 sundae is 1.2 so earning per min = 1.2 *2 F(t) = 1.2*2*t